# Transfer Learning in ECG Diagnosis: Is It Effective?

### v1

@@ 1 Introduction
Electrocardiogram (ECG) signals play a critical role in the early detection and diagnosis of cardiovascular diseases. The integration of automatic ECG interpretation, fueled by digitization and deep learning, has demonstrated performance on par with cardiologists [1, 2]. A major challenge to wide-scale adaptation of deep learning to ECG diagnosis is the lack of large-scale, high-quality labeled datasets in most real-world scenarios, due to prohibitive collection and annotation costs. To overcome this challenge, transfer learning is commonly employed, wherein features and parameters learned from a large dataset are reused and fine-tuned on a typically smaller, new dataset. This technique has been adapted from computer vision [3, 4, 5, 6, 7] to the ECG domain. Some studies have applied transfer learning to classify ECG arrhythmia by borrowing pre-trained weights on 2-D ImageNet [8], after transforming 1-D ECG signals to 2-D representations. For example, Salem et al. [9] generated 2-D spectrograms from ECG using Fourier Transform and applied pre-trained weights of DenseNet [10] on ImageNet [8] to detect ventricular fibrillation, atrial fibrillation, and ST-changes. Tadesse et al. [11] also employed spectrograms and leveraged pre-trained inception-v3 GoogLeNet [12] to diagnose cardiovascular diseases. Gajendran et al. [13] and Venton et al. [14] leveraged scalogram for 2-D conversion and reusing convolutional neural networks (CNNs) [10, 12, 15, 16, 17, 18] pre-trained on ImageNet to classify ECG records. Zhang et al. [19] applied Hilbert Transform and Wigner-Ville distribution [20, 21] to convert signals to 2-D then using pre-trained ResNet101 [16] to build their classifiers.
@@
Additionally, applying transfer learning directly to 1-D signals has shown encouraging results. Strodthoff et al. [26] reported significant improvements when pre-training xresnet1d101 [27] on the PTB-XL [25] dataset and subsequently fine-tuning on smaller datasets. Weimann et al. [28] achieved up to a 6.57% improvement in the classification performance of Atrial Fibrillation using CNNs, pre-trained on the large Icentia11K dataset [29]. Jang et al. [30] showed that pre-training a convolutional autoencoder on the AUMC ICU dataset [31] of size 26,481 worked better than training from random initialization on the 10,646-sample dataset of the Shaoxing People’s Hospital of China [32]. Other studies have also reported positive results of transfer learning [33, 34, 35, 36, 37].
While previous 1-D approaches have demonstrated the effectiveness of transfer learning in ECG diagnosis, these studies often focused on specific datasets and model architectures. An implicit assumption is that transferring knowledge from a large upstream dataset consistently improves downstream performance on another dataset, compared to training from random initialization (scratch). However, this hypothesis has not been systematically verified. In this study, we aim to validate the hypothesis by testing it across different ECG datasets and deep learning architectures. Specifically, we conduct extensive experiments using three upstream datasets for pre-training models and five downstream datasets for fine-tuning pre-trained models. We employ six deep learning models, encompassing the two predominant architectures for ECG diagnosis: Convolutional Neural Networks and Recurrent Neural Networks (RNNs). The comparison between fine-tuning performance and training from scratch provides insights into the effectiveness of transfer learning in ECG applications. Our key contributions and findings are as follows:
@@ 2.1 Datasets
Refer to caption (a) Pre-trained on PTB-XL
Refer to caption (b) Pre-trained on CPSC2018
Refer to caption (c) Pre-trained on Georgia
Refer to caption (a) Pre-trained on PTB-XL
Refer to caption (b) Pre-trained on CPSC2018
Refer to caption (c) Pre-trained on Georgia
@@ 2.2.2 Pre-training
[…]
@@ 3.1 Fine-tuning does not necessarily improve performance
3.2 Fine-tuning improvement fade with downstream dataset size
@@ 4 Conclusion
Moreover, our results confirm that fine-tuning tends to yield more effective results with CNNs than with RNNs in ECG classification. Contrary to 2-D images, RNNs are also a potential method to process time-series ECG signals. However, as mentioned in Section 3.4, inherent designs of RNNs make it more difficult to transfer knowledge learned from one dataset to another. Even in the case of training from scratch, LSTM, Bi-LSTM, and GRU showed inferior performance than that of ResNets (Section LABEL:compare_finetune). Thus we argue that in general, CNNs should be the preferred choice when deciding on architectures for ECG applications.
@@ Acknowledgement
This section provides the full numerical results of our experiments in Section LABEL:compare_finetune and 3.4.

### v2

@@ 1 Introduction
Electrocardiogram (ECG) signals play a critical role in the early detection and diagnosis of cardiovascular diseases. The integration of automatic ECG interpretation, fueled by digitization and deep learning, has demonstrated performance on par with cardiologists [1, 2]. A major challenge to wide-scale adaptation of deep learning to ECG diagnosis is the lack of large-scale, high-quality labeled datasets in most real-world scenarios, due to prohibitive collection and annotation costs. To overcome this challenge, transfer learning is commonly employed, where features and parameters learned from a large dataset are reused and fine-tuned on a typically smaller, new dataset. This technique has been adapted from computer vision [3, 4, 5, 6, 7] to the ECG domain. Some studies have applied transfer learning to classify ECG arrhythmia by borrowing pre-trained weights on 2-D ImageNet [8], after transforming 1-D ECG signals to 2-D representations. For example, Salem et al. [9] generated 2-D spectrograms from ECG using Fourier Transform and applied pre-trained weights of DenseNet [10] on ImageNet [8] to detect ventricular fibrillation, atrial fibrillation, and ST-changes. Tadesse et al. [11] also employed spectrograms and leveraged pre-trained inception-v3 GoogLeNet [12] to diagnose cardiovascular diseases. Gajendran et al. [13] and Venton et al. [14] leveraged scalogram for 2-D conversion and reusing convolutional neural networks (CNNs) [10, 12, 15, 16, 17, 18] pre-trained on ImageNet to classify ECG records. Zhang et al. [19] applied Hilbert Transform and Wigner-Ville distribution [20, 21] to convert signals to 2-D then using pre-trained ResNet101 [16] to build their classifiers.
@@
Additionally, applying transfer learning directly to 1-D signals has shown encouraging results. Strodthoff et al. [26] reported significant improvements when pre-training xresnet1d101 [27] on the PTB-XL [25] dataset and subsequently fine-tuning on smaller datasets. Weimann et al. [28] achieved up to a 6.57% improvement in the classification performance of Atrial Fibrillation using CNNs, pre-trained on the large Icentia11K dataset [29]. Jang et al. [30] showed that pre-training a convolutional autoencoder on the AUMC ICU dataset [31] of size 26,481 worked better than training from random initialization on the 10,646-sample dataset of the Shaoxing People’s Hospital of China [32]. Other studies have also reported positive results of transfer learning [33, 34, 35, 36, 37, 38].
While previous 1-D approaches have demonstrated the effectiveness of transfer learning in ECG diagnosis, these studies often focused on specific datasets and model architectures. An implicit assumption is that transferring knowledge from a large upstream dataset consistently improves downstream performance on another dataset, compared to training from random initialization (scratch). However, this hypothesis has not been systematically verified. In this study, we aim to validate the hypothesis by testing it across different ECG datasets and deep learning architectures. Specifically, we conduct extensive experiments using three upstream datasets for pre-training models and five downstream datasets for fine-tuning pre-trained models. We employ six deep learning models, encompassing the two predominant architectures for ECG diagnosis: Convolutional Neural Networks [39, 40, 41, 42, 43, 44, 45, 46, 47] and Recurrent Neural Networks (RNNs) [48, 49, 50, 51, 46, 43, 52, 53]. The comparison between fine-tuning performance and training from scratch provides insights into the effectiveness of transfer learning in ECG applications. Our key contributions and findings are as follows:
@@ 2.1 Datasets
(a) Pre-trained on PTB-XL
(b) Pre-trained on CPSC2018
(c) Pre-trained on Georgia
@@ 2.2.2 Pre-training
In this work, we examined six DNN architectures. Three of these were convolutional: ResNet1d18, ResNet1d50, and ResNet1d101, which were adapted from the original 2-D versions [16]. The other three were recurrent DNNs: Long Short Term Memory (LSTM) [55], Bidirectional LSTM [55], and Gated Recurrent Unit (GRU) [56]. Three datasets were used for pre-training: PTB-XL, CPSC2018, and Georgia, due to their substantial sample sizes. Each of the six models was pre-trained on the training subset of each dataset for 100 epochs using Adam optimizer [57] with learning rate $0.01$. We evaluated each model on the test subset during training, and only the checkpoint that achieved the best evaluation metric over 100 epochs was saved as the pre-trained model. We opted not to save the last checkpoint at the 100th epoch due to observed overfitting as training progressed, especially for the three recurrent models.
@@ 3.1 Fine-tuning does not necessarily improve performance
3.2 Fine-tuning improvement fades with downstream dataset size
@@ 4 Conclusion
Moreover, our results confirm that fine-tuning tends to yield more effective results with CNNs than with RNNs in ECG classification. Contrary to 2-D images, RNNs are also a potential method to process time-series ECG signals. However, as mentioned in Section 3.4, inherent designs of RNNs make it more difficult to transfer knowledge learned from one dataset to another. Even in the case of training from scratch, LSTM, Bi-LSTM, and GRU showed inferior performance than that of ResNets (Section 3.1). Thus we argue that in general, CNNs should be the preferred choice when deciding on architectures for ECG applications.
@@ Acknowledgement
This section provides the full numerical results of our experiments in Section 3.1 and 3.4.